# M. C. Rodriguez

### v3

@@ Abstract
We will study in details the full scalar potential of the Minimal Supersymmetric $SU(3)_{C}\times SU(3)_{L}\times U(1)_{N}$ Model when all the scalars fields belong to sextet and anti-sextet representation obtain VEV. We will present numerical predictions for neutral scalars of this model, we will show that the mass of the light scalar of CP par is $125.5$ GeV while the light pseudoscalar have mass bigger than $150$ GeV.
@@ 1 Introduction
$\phi$ $=$ $\left(\begin{array}[]{c}h^{+}\\ h^{0}\end{array}\right)\sim({\bf 1},{\bf 2},+1).$ (6)
@@
$h^{0}=\frac{1}{\sqrt{2}}\left(v+\phi_{1}\pm\imath\phi_{2}\right),$ (12)
under this expansion, our scalar potential become
$V(\phi)=V(h^{0})-2\mu^{2}\phi^{2}_{1}+hv\phi_{1}\left(\phi^{2}_{1}+\phi^{2}_{2}\right)+\frac{h}{4}\left(\phi^{2}_{1}+\phi^{2}_{2}\right)^{2},$ (13)
where $\phi_{1}$ is CP even field and it get the following mass [^2] A nice review about it can be found in [12].
@@
we must emphasize the SM do not predict any value for this mass, in fact it can get any values rom 00 until any finite values, but we can get a superior limit in its mass given by $M_{H}<1000$ GeV [10, 11]. By another hand, the field $\phi_{2}$ is CP odd field and it as massless field (Goldstone particles) and will be the responsable to give mass for $Z$ gauge boson, this mechanism is know as Brout-Englert-Higgs mechanism [12, 13]. The mass of the light Higgs boson $h^{0}$ was measured at the LHC and its value is given by [14, 15]
$M_{H}=125.09\pm 0.21\,\mathrm{(stat.)}\pm 0.11\,\mathrm{(syst.)}~{}\mathrm{GeV}.$ (15)
@@
it is know as Gelmini-Roncadelli scheme [24]. The mass generation mechanism of charged leptons remains the same as presented in the SM.
@@
Figure 2: We show the predections for the parameter $\rho$, see our Eq.(22), in the scheme of Gelmini-Roncadelli as a function of the Triplet VEV $v_{\Delta}$.
@@
the mass $m_{h}$ as function of the VEV $v_{\Sigma}$ is shown in our Fig.(3), where we can explain the experimental. For the case of CP odd sector the mass matrix has the following representation
@@
Figure 3: We show the predections for the light higgs mass $h$ in the scheme of Gelmini-Roncadelli as a function of the parameter $\lambda_{1}$ and we show that we can explain the data from Eq.(15).
@@
where $V_{G}(\phi,\Delta)$ is defined in our Eq.(LABEL:potmajorongelmini) and $\tilde{\phi}\equiv\left(\imath\sigma_{2}\phi\right)$. Under this hipotesis our Eq.(47) become
@@
$m_{H}$ as function of the VEV $v_{\Delta}$ is shown in our Fig.(4).
Figure 4: We show the predections for the heavy Higgs mass $H$ in the scheme of Gelmini-Roncadelli, see Eq.(48), as a function of the Triplet Field VEV $V_{\Delta}$.
@@
$m_{M}$ as function of the VEV $v_{\Delta}$ is shown in our Fig.(5).
Figure 5: We show the predections for the Majoron mass $M$, see Eq.(50), in the scheme of Ma as a function of the Triplet Field VEV $V_{\Delta}$.
@@
Another interesting possibility is to build Supersymmetric models with Higgs triplets as done in [45, 46, 47, 48]. There is also, an interesting proposal to explain this anomaly using the Minimal $R$-Symmetric extension of the Minimal Superymmetric Standard Model, known as MRSSM [49, 50, 51], and in this case we must respect the following inequality $|v_{\Delta}|\leq 4$ GeV [50].
@@ 3 Case 1
coming from $M_{W}$, where, we have defined $V^{2}_{\eta}=v^{2}_{\eta}+v^{\prime 2}_{\eta}$ and $V^{2}_{\rho}=v^{2}_{\rho}+v^{\prime 2}_{\rho}$ and $V^{2}_{2}=v^{2}_{\sigma_{2}}+v^{2}_{\sigma^{\prime}_{2}}$. Assuming that $v_{\eta}=20$, $v_{\chi}=1500$, $v_{\sigma_{2}}=10$, $v^{\prime}_{\eta}=v^{\prime}_{\rho}=v^{\prime}_{\sigma_{2}}=v^{\prime}_{\chi}=1$ in GeV, the value of $v_{\rho}$ is fixed by the constraint above. We get similar values are presented in [62]. We present in our Fig.(6) our results for the masses of $M_{h}$ and $M_{A}$ and we see it is in agreement with Eq.(15).
Figure 6: The masses of light scalars, $M_{h}$ and light pseudoscalar $M_{A}$, in terms of $v_{\chi}$ for $v_{\chi^{\prime}}=10$ GeV and $M^{exp}_{h}$ is the experimental values given at our Eq.(15).
@@ 4 Case 2
Our results are shown in Figs.(7,8,9,10) where we can say
@@
However if we take $f_{2}$, $f_{4}$, $f^{\prime}_{2}$ and $f^{\prime}_{4}$ not no null the Majoron dissapear but it has mass of few GeV, the same results as presented in [25], but the values of $M_{A}$ is near the ones we present in our Figs.(7,9).
@@ 5 Conclusions
We have studied the scalar potential of supersymmetric $331$ model with all the VEV from the sextet and anti-sextet not no vanishing and we shown our light scalar fields has mass near the value $M_{h}\approx 125.5$ GeV, with concordance with the experimental values for Higgs masses given by Eq.(15), and $M_{A}>150\,\ {\mbox{G}eV}$.

### v4

@@ Abstract
We will study in details the full scalar potential of the Minimal Supersymmetric $SU(3)_{C}\times SU(3)_{L}\times U(1)_{N}$ Model when all the scalars fields belong to sextet and anti-sextet representation obtain VEV. We will present numerical predictions for neutral scalars of this model, we will show that the mass of the light scalar of CP par is $125.5$ GeV while the light pseudoscalar have mass bigger than $400$ GeV.
@@ 1 Introduction
$\phi=\left(\begin{array}[]{c}\phi^{+}\\ \phi^{0}\end{array}\right)\sim({\bf 1},{\bf 2},+1).$ (6)
@@
$\phi^{0}=\frac{1}{\sqrt{2}}\left(v+h^{0}+\imath G^{0}\right),$ (12)
under this expansion, our scalar potential become [10]
$V(\phi)$ $=$ $V(\phi^{\prime 0})-\mu^{2}\left(h^{0}\right)^{2}+\lambda vh^{0}\left((h^{0})^{2}+(G^{0})^{2}\right)+\frac{\lambda}{4}\left((h^{0})^{2}+(G^{0})^{2}\right)^{2},$
$V(\phi^{\prime 0})$ $=$ $-\frac{\lambda}{4}v^{4},\,\ |\phi^{\prime 0}|=\sqrt{\frac{-\mu^{2}}{2\lambda}}\equiv\frac{v}{\sqrt{2}}>0,$ (13)
where $h^{0}$ is CP even field and it get the following mass [^2] A nice review about this interesting subject can be found in a very nice review [11].
@@
we must emphasize the SM do not predict any value for this mass, in fact it can get any values from 00 until any finite values, but we can get a superior limit given by $M_{H}<1000$ GeV [12, 13]. By another hand, the field $G^{0}$ is CP odd field and it as massless field (Goldstone particles) and will be the responsable to give mass for $Z$ gauge boson, this mechanism is know as Brout-Englert-Higgs mechanism [11, 14]. The mass of the light Higgs boson $h^{0}$ was measured at the LHC and its value is given by [15, 16]
$M_{H}=125.20\pm 0.21\,\mathrm{(stat.)}\pm 0.11~{}\mathrm{GeV}.$ (15)
@@
it is know as Gelmini-Roncadelli scheme [25]. The double beta decay experiments can show if the neutrinos are Majorna particle or not [26, 27, 28, 29] The mass generation mechanism of charged leptons remains the same as presented in the SM.
@@
Figure 2: We show the predections for the parameter $\rho$, see our Eq.(22), in the scheme of Gelmini-Roncadelli as a function of the Triplet VEV $v_{\Delta}$. $\rho^{cen}=0.9998$ while $\rho^{inf}=\rho^{cen}-0.0008$
@@
the mass $M_{h}$ as function of the parameter $\lambda_{1}$ is shown in our Fig.(3), where we can explain the experimental given by Eq.(15). For the case of CP odd sector the mass matrix has the following representation
@@
Figure 3: We show the predections for the light higgs mass $h$ in the scheme of Gelmini-Roncadelli as a function of the parameter $\lambda_{1}$ and $M^{exp}_{h}$ is the central experimental values given at our Eq.(15) , but we will present those in more detail below. $M^{sup}_{h}=M^{exp}_{h}+0.11$ while $M^{inf}_{h}=M^{exp}_{h}-0.11$.
@@
where $V_{G}(\phi,\Delta)$ is defined in our Eq.(LABEL:potmajorongelmini) and $\tilde{\phi}\equiv\left(\imath\sigma_{2}\right)\phi$. Under this hipotesis our Eq.(47) become
@@
$M_{H}$ as function of the VEV $v_{\Delta}$ is shown in our Fig.(4). The data for the heavy Higgs sector as well for the pseudoscalar $A$ is [74]
[…]
GeV.
Figure 4: We show the predections for the heavy Higgs mass $H$ in the scheme of Gelmini-Roncadelli, see Eq.(48) and $\mu=1$ GeV, as a function of the Triplet Field VEV $V_{\Delta}$ and $M^{exp}_{H}$ is the lower experimental values given at our Eq.(49).
@@
$m_{M}$ as function of the VEV $v_{\Delta}$ is shown in our Fig.(5) and our results for the mass of $M_{M}$ is in agreement with Eq.(49), while for the case of $u=0.1$ GeV and $\mu=1$ GeV we get $M_{H}=925.11$ GeV and $M_{M}=777.92$ GeV and those values are in agreement with the experimental data giving in our Eq.(49) and to be in agreement with the exerimental data we must have
$\mu>10^{-1}$ (52)
in GeV. For the case of $\mu=10^{(}14)$ GeV the results are shown in our Figs.(6,7).
Figure 5: We show the predections for the Majoron mass $M$, see Eq.(51) and $\mu=1$ GeV, in the scheme of Ma as a function of the Triplet Field VEV $V_{\Delta}$ and $M^{exp}_{A}$ is the lower experimental values given at our Eq.(49).
Figure 6: We show the predections for the heavy Higgs mass $H$ in the scheme of Gelmini-Roncadelli, see Eq.(48) and $\mu=10^{14}$ GeV, as a function of the Triplet Field VEV $V_{\Delta}$.
Figure 7: We show the predections for the Majoron mass $M$, see Eq.(51) and $\mu=10^{14}$ GeV, in the scheme of Ma as a function of the Triplet Field VEV $V_{\Delta}$.
@@
There are also an experimental version of the Gelmini-Rocandeli scheme [50, 51] where as in the MSSM we have two Higgs in the doublet representations
$H_{1}$ $=$ $\left(\begin{array}[]{c}h^{0}_{1}\\ h^{-}_{1}\end{array}\right)\sim({\bf 1},{\bf 2}^{*},-1),\,\ <H_{1}>=\frac{1}{\sqrt{2}}\left(\begin{array}[]{c}v_{1}\\ 0\end{array}\right),$ (57)
$H_{2}$ $=$ $\left(\begin{array}[]{c}h^{+}_{2}\\ h^{0}_{2}\end{array}\right)\sim({\bf 1},{\bf 2},+1),\,\ <H_{2}>=\frac{1}{\sqrt{2}}\left(\begin{array}[]{c}0\\ v_{2}\end{array}\right),$ (62)
besides we also add one scalar in the triplet representation ${\bf T}_{1}=(T_{1},T_{2},T_{3})$ and another one in the anti-triplet representation ${\bf T}_{2}=(\bar{T}_{1},\bar{T}_{2},\bar{T}_{3})$, where
$\Delta_{1}$ $=$ $\left(\imath\sigma_{2}\right){\bf T}_{1}=\left(\begin{array}[]{cc}\Delta^{0}_{1}&\frac{h^{+}_{1}}{\sqrt{2}}\\ \frac{h^{+}_{1}}{\sqrt{2}}&-H^{++}_{1}\end{array}\right)\sim\left({\bf 1},{\bf 3},2\right),\,\ \langle\Delta_{1}\rangle=\frac{1}{\sqrt{2}}\left(\begin{array}[]{cc}v_{\Delta_{1}}&0\\ 0&0\end{array}\right),$ (67)
$\Delta_{2}$ $=$ $\left(\imath\sigma_{2}\right){\bf T}_{2}=\left(\begin{array}[]{cc}\Delta^{--}_{2}&\frac{h^{-}_{2}}{\sqrt{2}}\\ \frac{h^{-}_{2}}{\sqrt{2}}&-\Delta^{0}_{2}\end{array}\right)\sim\left({\bf 1},{\bf 3}^{3},-2\right),\,\ \langle\Delta_{2}\rangle=\frac{1}{\sqrt{2}}\left(\begin{array}[]{cc}0&0\\ 0&v_{\Delta_{2}}\end{array}\right).$ (72)
In the supersymmetric version of Gelmini-Rocandeli model we introduce all the fields in the superchiral superfields and the superpotential of this model is [50, 51]
$W^{GR}_{SUSY}$ $=$ $W^{MSSM}_{RC}+W^{MSSM}_{RHN}+\mu_{\Delta}\hat{\Delta}_{1}\hat{\Delta}_{2}+\frac{f^{N}_{ij}}{2}\left(\hat{L}_{i}\hat{L}_{j}\right)\hat{\Delta}_{1}+\lambda_{1}\hat{H}_{1}\Delta_{1}\hat{H}_{1}$
$+$ $\lambda_{2}\hat{H}_{2}\Delta_{2}\hat{H}_{2},$
$W^{MSSM}_{RC}$ $=$ $\mu\hat{H}_{1}\hat{H}_{2}+f^{l}_{ij}\left(\hat{H}_{1}\hat{L}_{i}\right)\hat{E}_{j}+f^{d}_{ij}\left(\hat{H}_{1}\hat{Q}_{i}\right)\hat{D}_{j}+f^{u}_{ij}\left(\hat{H}_{2}\hat{Q}_{i}\right)\hat{U}_{j},$
$W^{MSSM}_{RHN}$ $=$ $\frac{M_{ij}}{2}\hat{N}_{i}\hat{N}_{j}+\frac{f^{\nu}_{ij}}{2}\left(\hat{H}_{2}\hat{L}_{i}\right)\hat{N}_{j},$
where $W^{MSSM}_{RC}$ is the tradition superpotential of MSSM with $R$ parity conservation and you can see it in [52] while $W^{MSSM}_{RHN}$ is the superpotential of three right-handed neutrinos [53]. Therefore the mass mechanism for the neutrinos is the same as in the Gelmini-Rocandeli model and is presented in [50]. The mass of the lighest higgs $h^{0}$ is still gave by our Eq.(44) and for the heavy higgs $H$ by our Eq.(48) and for the Majoron $M$ it is given at Eq.(51).
The masses of the gauge bosons in this model are
$M^{2}_{W}$ $=$ $\frac{g^{2}}{4}\left[v^{2}+2\left(v^{2}_{\Delta_{1}}+v^{2}_{\Delta_{2}}\right)\right],$
$M^{2}_{Z}$ $=$ $\frac{g^{2}}{4}\left[v^{2}+4\left(v^{2}_{\Delta_{1}}+v^{2}_{\Delta_{2}}\right)\right],$ (75)
and we obtain the same results that we presented in our study [33]. The usual scalar sector of this model is presented in [54, 55].
@@ 3 Case 1
coming from $M_{W}$, where, we have defined $V^{2}_{\eta}=v^{2}_{\eta}+v^{\prime 2}_{\eta}$ and $V^{2}_{\rho}=v^{2}_{\rho}+v^{\prime 2}_{\rho}$ and $V^{2}_{2}=v^{2}_{\sigma_{2}}+v^{2}_{\sigma^{\prime}_{2}}$. Assuming that $v_{\eta}=20$, $v_{\chi}=1500$, $v_{\sigma_{2}}=10$, $v^{\prime}_{\eta}=v^{\prime}_{\rho}=v^{\prime}_{\sigma_{2}}=v^{\prime}_{\chi}=1$ in GeV, the value of $v_{\rho}$ is fixed by the constraint above. We get similar values are presented in [66]. We present in our results in our Figs.(8,9) and Fig.(9) our results for the masses of $M_{h}$, $M_{H}$ and $M_{A}$ and we see it is in agreement with Eqs.(15,49) and it is easy to conclude the following results
$900\leq v_{\chi}\leq 1000,\,\ M_{h}\approx 125;$
$v_{\chi}>1800,\,\ M_{H}\approx M_{A}>400$ (128)
Figure 8: The masses of light scalars, $M_{h}$, in terms of $v_{\chi}$ for $v_{\chi^{\prime}}=1$ GeV and $M^{exp}_{h}$ is the central experimental values given at our Eq.(15) , but we will present those in more detail below. $M^{sup}_{h}=M^{exp}_{h}+0.11$ while $M^{inf}_{h}=M^{exp}_{h}-0.11$.
Figure 9: The masses of the usual heavy scalar, $M_{H}$, in terms of $v_{\chi}$ for $v_{\chi^{\prime}}=1$ GeV and $M^{exp}_{H}$ is the lower experimental values given at our Eq.(49).
Figure 10: The masses of the light pseudoscalar, $M_{A}$, in terms of $v_{\chi}$ for $v_{\chi^{\prime}}=1$ GeV and $M^{exp}_{A}$ is the lower experimental values given at our Eq.(49).
@@ 4 Case 2
Our results are shown in Figs.(11,12,13,14) where we can say
@@
However if we take $f_{2}$, $f_{4}$, $f^{\prime}_{2}$ and $f^{\prime}_{4}$ not no null the Majoron dissapear but it has mass of few GeV, the same results as presented in [30], but the values of $M_{A}$ is near the ones we present in our Figs.(11,13).
@@ 5 Conclusions
We have studied the scalar potential of supersymmetric $331$ model with all the VEV from the sextet and anti-sextet not no vanishing and we shown our light scalar fields has mass near the value $M_{h}\approx 125.5$ GeV, with concordance with the experimental values for Higgs masses given by Eq.(15), and $M_{A}>400\,\ {\mbox{G}eV}$.